Monique Diamond-Barnes inherited $350 million from her late husband Biff Barnes and has decided to use some of the money to purchase an NBA team so that she can use the team to market her line of expensive perfumes. She is considering purchasing a team in Sacremento for an asking price of $300 million. As her accountant, Monique has asked you to estimate her potential profit from the club over the first five years of operation. You anticipate that annual revenues will be $200 million and annual operating costs will include $95 million in player expenses, $30 million in administrative expenses and $35 million in rental for the arena. The corporate tax rate on net income is 40%

a) Compute the annual expected net income for the team before and after taxes. What is the annual rate of return as a percentage of the purchase price?

b) Monique doesn’t think that the team will earn quite enough profit for her flailing perfume business.
Can you think of a way to increase her after-tax annual rate of return? If so, compute it.

c) At the end of the first five years of operations, Monique is growing bored of being a basketball team owner,
  but might be convinced to have her spoilt son take over the team operations with Monique still retaining ownership.
  What are the options open to her son to do with the team?.
Answer #1

INCOME EARNED ANNUALLY

Revenue: 200 million

(-) expenses:

*(95+35+30). (160) million

Revenue before tax=40 million

(-) tax@40%=16million

Net annual revenue=24million

For 5 year=5*24=120million{no profit till end of 5 years of investment of 400 million}

#annual rate of return=annual revenue÷purchase price

=24÷300=8%

#). Her son needs to carry the Team more years of 7.5 years